Find the expectation of a random variable X if f(x) = ce⁻ ˣ for x>0 and 0 otherwise.


1
Expert's answer
2021-09-06T16:16:58-0400

Solution:

"\int_0^{\infty} f(x)dx=1
\\\Rightarrow\int_0^{\infty} (ce^{-x})dx=1
\\\Rightarrow -c[e^{-x}]_0^{\infty}=1
\\\Rightarrow -c[0-1]=1
\\\Rightarrow c=1"

Then,

"E[X]=\int_0^{\infty}x f(x)dx
\\=\int_0^{\infty}x (ce^{-x})dx
\\=\int_0^{\infty}x e^{-x}dx
\\=[-e^{-x}x-e^{-x}]_0^{\infty}
\\=[0-0]-[0-1]
\\=1"